# Time-Aware One Step Diffusion Network for Real-World Image Super-Resolution

### v2

@@ Introduction
As shown in Fig. 3, with the same input image, TAE encodes it into different latent feature condition on the timestep $t_{s}$. Overall, as the $t_{s}$ increases, both the mean and variance of the latent features show a decreasing trend. After visualizing the latent feature via PCA dimensionality reduction, we can also clearly observe the changes in the latent space.
Time-Aware Variational Score Distillation
Following the OSEDiff (Wu et al. 2024b), the VSD loss has been widely adopted in SD-based one-step Real-ISR methods to enhance the realism of reconstruction results. However, these methods generally overlook that the guidance provided by VSD at different timesteps is actually inconsistent. As shown in Fig. 4, the mean and standard deviation of the VSD loss exhibit a clear upward trend as the timestep increases. When decoding the results at different timesteps into the pixel space, we observe that at $t_{v}$ equals 100, the outputs of the teacher and LoRA models are similar, and the gradients mainly reflect enhancements in texture details. In contrast, when $t_{v}$ increases to 300, the teacher model’s output contains significantly more semantic information while the LoRA model’s output remains smooth, and the gradients reflect global semantic guidance. However, when $t_{v}$ increases to 600, the teacher model can only recover coarse color and structural information from the noisy latent input, making it difficult to provide meaningful guidance. This implies that the VSD loss provides distinct guidance for the same image depending on $t_{v}$. Such opposing directional guidance creates conflicting optimization signals for the student model, ultimately leading to suboptimal convergence.
@@
DRealSR
PSNR $\uparrow$ 28.030 25.929 28.073 28.345 27.539 27.915 28.318 26.20 28.099 28.387
@@
RealSR
PSNR $\uparrow$ 24.645 24.240 25.149 26.266 25.183 25.148 25.503 23.404 25.469 25.166
SSIM $\uparrow$ 0.7080 0.6649 0.7211 0.7341 0.7269 0.7338 0.7418 6886 0.7301 0.7150
@@
RealLR200
CLIPIQA $\uparrow$ 0.6036 0.7072 0.7023 0.6474 0.7122 0.6792 0.7153 0.7248 0.7048 0.7741
@@
This phenomenon arises because most of the image information is preserved when $t$ is small, and both the teacher and LoRA models focus mainly on generating texture details. As $t$ increases, the injection of more noise gradually obscures the underlying image content, forcing the teacher model to rely more heavily on its generative prior. However, since the LoRA model is trained on low-quality data generated by the student model and does not employ the CFG strategy (Ho and Salimans 2022), its outputs tend to be overly smooth.
@@
We then add noise to $\hat{z}$ corresponding to $t_{v}$ to obtain $\hat{z}_{t_{v}}=\alpha_{t_{v}}\hat{z}+\beta_{t_{v}}\epsilon$, which is then passed through the teacher model and the LoRA model with $t_{v}$ to compute the TAVSD loss:
@@
Training Loss
We train the student model with reconstruction and regression losses. To avoid gradient inconsistency arising from the ill-posed problem of the Real-ISR task (Liang, Zeng, and Zhang 2022) while fully leveraging the teacher model knowledge introduced by VASVD, we first apply a Gaussian blur to both the reconstructed image and the HQ image $x_{H}$ before computing the MSE loss. This ensures that the $x_{H}$ only supervises the low-frequency content of the reconstruction, helping to preserve high-frequency details. We adopt a larger blur kernel for larger timesteps $t_{s}$, which enhances the trade-off between fidelity and generation:
@@
Where * denotes the convolution operation, $G_{t_{s}}$ is the convolution kernel whose size is determined by $t_{s}$. This loss and the LPIPS loss, forms the reconstruction loss:
@@
For the regression loss, we adopt the TAVSD loss in Eq. 5 to improve the realism of the generated results. The overall loss for the student model is:
@@
For the LoRA model, we adopt the original diffusion loss:
@@
where $\epsilon^{\prime}$ is the Gaussian noise as the training target for the denoising network.
Experiments
Experimental Setup
@@
Training. We use LSDIR (Li et al. 2023) as the training data with the $512\times 512$ patch size. To generate paired HQ-LQ training data, we follow the degradation pipeline from Real-ESRGAN (Wang et al. 2021). We use AdamW optimizer (Loshchilov and Hutter 2017) with a learning rate $5\times 10^{-5}$ and set LoRA rank to 4 for both the student model and LoRA model. We employ the SD 2.1-base as the pre-trained diffusion model and fine-tune it for 2k iterations using 8 NVIDIA A40 GPUs with a batch size of 24.
@@
Test Dataset. We evaluate our method in both synthetic and real-world dataset. For the synthetic dataset, we randomly crop 3K patches with a resolution of 512 × 512 from the DIV2K (Agustsson and Timofte 2017) validation set and synthesize LQ data using the same pipeline as that in training. For real-world data, we employ RealSR (Cai et al. 2019), DrealSR (Wei et al. 2020), and RealLR200 (Wu et al. 2024c). We center-crop RealSR (Cai et al. 2019) and DrealSR (Wei et al. 2020) datasets with size $128\times 128$ for LQ images and $512\times 512$ for HQ image. For RealLR200 (Wu et al. 2024c) dataset, since the corresponding HQ images are unavailable, we perform only a $128\times 128$ center-crop on the LQ images.
@@
Evaluation Metrics. We utilize several reference and non-reference metrics to evaluate the performance of various methods on the test data. For the reference measures, we employ PSNR, SSIM (Wang et al. 2004), and LPIPS (Zhang et al. 2018a) to measure image fidelity. For the non-reference measures, we employ CLIPIQA (Wang, Chan, and Loy 2023), MUSIQ ((Ke et al. 2021), MANIQA (Yang et al. 2022), TOPIQ (Chen et al. 2024), and QALIGN (Wu et al. 2024a) to measure image quality.
@@
Compared Methods. We compare our method with several multi-step diffusion-based methods StableSR (Wang et al. 2023a), DiffBIR (Lin et al. 2023), SeeSR (Wu et al. 2024c)), and one-step diffusion-based methods SinSR (Wang et al. 2024), OSEDiff (Wu et al. 2024b), S3Diff (Zhang et al. 2024), AdcSR (Chen et al. 2025), TSDSR (Dong et al. 2025), and PisaSR (Sun et al. 2025). All comparative results are obtained using publicly released code for testing.
Comparisons with State-of-the-art Methods
@@
Quantitative Comparisons. We set up the timestep condition $t_{s}=500$ in our method, and show the quantitative comparisons on the four synthetic and real-world datasets in Tab. 1. We have the following observations: (1) TADSR achieves the highest no-reference scores across four datasets, except for the MUSIQ on DIV2K-Val. This demonstrates that TADSR can more effectively leverage the generative priors from SD to produce more realistic results. Notably, TADSR is the only one-step methods that consistently outperforms multi-step methods on all no-reference metrics, achieving both efficiency and perceptual quality. (2) TADSR maintains PSNR values comparable to other SD-based one-step methods, indicating a good balance between fidelity and realism. (3) TADSR shows clear improvements over other SD-based one-step methods on CLIPIQA and TOPIQ, highlighting its superior semantic awareness and generative capability.
Figure 6: (a) Quantitative metrics of our method under different timestep condition $t_{s}$, evaluated on the DrealSR dataset. (b) Comparison of our method under different timesteps $t_{s}$, PisaSR under different semantic guidance weights $\lambda_{sem}$, and other one-step diffusion-based Real-ISR methods, evaluated on the DrealSR dataset.
@@
Qualitative Comparisons. Fig. 5 shows the visual comparisons between our method and the other state-of-the-art Real-ISR methods. As shown in the first row, TADSR generates significantly more natural and sharper textures from heavily degraded LQ images, especially in facial regions such as the teeth, eyes, and eyebrows, demonstrating its strong semantic generation capability. In the second row, the digits and letters produced by TADSR appear much clearer, showcasing its superior degradation removal ability while preserving fidelity. In the third row, TADSR yields more natural results around the eagle’s eyes and beak. In the fourth row, only TADSR accurately restored natural-looking facial features such as the nose, mouth, and chin. Other methods generally suffered from degradation, resulting in some distortion, and failed to reconstruct a plausible chin structure. Overall, thanks to the ability to distill generative priors from SD more effectively in TAVSD loss, TADSR can produce natural and realistic results in a single diffusion step. Compared to other methods, it achieves strong perceptual quality while maintaining high efficiency.
Ablation Study
@@
Impact of Different Timestep Condition. As shown in Fig. 6(a), we analyze the impact of timestep $t_{s}$ in our method on both reference and no-reference metrics. As $t_{s}$ increases, PSNR exhibits a decreasing trend while QALIGN shows an upward trend, indicating a trade-off where fidelity is sacrificed to enhance realism. This trade-off between fidelity and realism aligns with the function of $t_{s}$, as a larger $t_{s}$ means that TAVSD provides stronger generative guidance, while a smaller $t_{s}$ provides more fidelity-preserving guidance. Similar visual results can be observed in supplementary materials. Furthermore, we compare the results of our method under different $t_{s}$, PisaSR under different $\lambda_{sem}$ settings, and other one-step Real-ISR methods, as shown in Fig. 6(b). It can be observed that our method consistently lies in the top-right corner across different $t_{s}$. When $t_{s}$ equals 200, our method achieves 26.61dB PSNR, which is more than 1dB higher than SinSR, and QALIGN is significantly higher than SinSR. In contrast, although PisaSR can also achieve a PSNR of 29.60dB by tuning the $\lambda_{pix}=1.0$ and $\lambda_{sem}=0.6$, its QALIGN is only 2.91, which is similar to SinSR. This indicates that our method achieves a substantial improvement in fidelity with only a minimal compromise in realism.
@@
Figure 7: Vision Comparisons of the ablation study on TAE and TAVSD. Baseline use the original VAE encoder in SD and VSD loss.
@@
Table 2: Quantitative Comparison of ablation study on TAVSD and TAE. Baseline uses the original VAE encoder in SD and VSD loss.
@@
Impact of TAVSD and TAE. To validate the effectiveness of TAVSD and TAE, we conducted ablation studies by removing them. We employ the original VAE encoder in SD and the VSD loss as our baseline and conduct ablation studies by separately removing TAE and TAVSD. We use PSNR to evaluate fidelity and MANIQA, MUSIQ, and TOPIQ to assess realism. As shown in Tab. 2, we have the following three key observations: (1) After removing TAE, both reference and no-reference metrics decline, demonstrating that adapting the latent feature according to timesteps helps more fully utilize the generative priors in SD. (2) When TAVSD is ablated, all metrics similarly decrease, indicating that more consistent guidance from the teacher model better activates generative priors across different timesteps. (3) Baseline shows significant degradation in PSNR and moderate decline in others, proving that both TAE and TAVSD improve fidelity and realism. Additionally, Fig. 7 presents a visual comparison of our ablation studies, showing that both the absence of TAE/TAVSD leads to unrealistic parrot reconstructions while the baseline even produces visible artifacts. In contrast, our method produces realistic and natural results by fully exploiting the generative priors in SD.
Conclusion

### v3

@@ 3.3 Time-Aware VAE Encoder
As shown in Figure 3, TAE encodes the same input image into different latent feature condition on the timestep $t_{s}$. Overall, as the $t_{s}$ increases, both the mean and variance of the latent features show a decreasing trend. After visualizing the latent feature via PCA dimensionality reduction, we can also clearly observe the changes in the latent space.
3.4 Time-Aware Variational Score Distillation
Following the OSEDiff [32], the VSD loss has been widely adopted in SD-based one-step Real-ISR methods to enhance the realism of reconstruction results. In one-step image generation tasks, the timestep is usually sampled randomly to distill the full generative prior of the teacher model. However, we find that in Real-ISR, this paradigm instead leads to inconsistent generative guidance, since SD exhibits different generative priors across timesteps, while the timestep sampling in the teacher model is completely random and independent of the student model.
As discussed in Section 3.1, due to the stop-gradient operation, the guidance of the VSD loss can be interpreted as the residual between the latent images predicted by the teacher model and the LoRA model. Therefore, we decode these latent images into the pixel space and analyze the guidance of the VSD loss at different timesteps. As illustrated in Figure 4, first, the mean and standard deviation of the VSD loss exhibit a clear upward trend as the timestep increases. Besides, we observe that at $t_{v}$ equals 100, the outputs of the teacher and LoRA models are similar, and the gradients mainly reflect enhancements in texture details. In contrast, when $t_{v}$ increases to 300, the teacher model’s output contains significantly more semantic information while the LoRA model’s output remains smooth, and the gradients reflect global semantic guidance. However, when $t_{v}$ increases to 600, the teacher model can only recover coarse color and structural information from the noisy latent input, making it difficult to provide meaningful guidance. This implies that the VSD loss provides distinct guidance for the same image depending on $t_{v}$. Such opposing directional guidance creates conflicting optimization signals for the student model, leading to suboptimal convergence.
@@ 3.4 Time-Aware Variational Score Distillation
DRealSR PSNR $\uparrow$ 28.030 25.929 28.073 28.345 27.539 27.915 28.318 26.20 28.099 28.387
@@
RealSR PSNR $\uparrow$ 24.645 24.240 25.149 26.266 25.183 25.148 25.503 23.404 25.469 25.166
SSIM $\uparrow$ 0.7080 0.6649 0.7211 0.7341 0.7269 0.7338 0.7418 0.6886 0.7301 0.7150
@@
RealLR200 CLIPIQA $\uparrow$ 0.6036 0.7072 0.7023 0.6474 0.7122 0.6792 0.7153 0.7248 0.7048 0.7741
@@
This phenomenon arises because most of the image information is preserved when $t$ is small, and both the teacher and LoRA models focus mainly on generating texture details. As $t$ increases, the injection of more noise gradually obscures the underlying image content, forcing the teacher model to rely more heavily on its generative prior. However, since the LoRA model is trained on low-quality data generated by the student model and does not employ the CFG strategy [11], its outputs tend to be overly smooth. This also explains why the VSD loss can make images more realistic while avoiding over-smoothing — since the LoRA model tends to produce smoother outputs than the teacher model, their residuals naturally provide high-frequency details that guide the generation process.
@@
We then add noise to $\hat{z}$ corresponding to $t_{v}$ to obtain $\hat{z}_{t_{v}}=\alpha_{t_{v}}\hat{z}+\beta_{t_{v}}\epsilon$, which is fed to the teacher model and the LoRA model with $t_{v}$ to compute the TAVSD loss:
@@
3.5 Training Loss
We train the student model with reconstruction and regression losses. To avoid gradient inconsistency arising from the ill-posed problem of the Real-ISR [15] while fully leveraging the teacher model knowledge, we first apply a Gaussian blur to both the reconstructed image and the HQ image $x_{H}$ before computing the MSE loss. This ensures that the $x_{H}$ only supervises the low-frequency content of the reconstruction, helping to preserve high-frequency details. We adopt a larger blur kernel for larger timesteps $t_{s}$, which enhances the trade-off between fidelity and generation:
@@ 3.5 Training Loss
Where * denotes the convolution operation, $G_{t_{s}}$ is the convolution kernel whose size is determined by $t_{s}$. This loss and the LPIPS loss form the reconstruction loss:
@@
For the regression loss, we adopt the TAVSD loss in Eq. (6) to improve the realism of the generated results. The overall loss for the student model is:
@@
We adopt the original diffusion loss for the LoRA model:
@@
where $\epsilon^{\prime}$ is the randomly sampled Gaussian noise as the training target for the denoising network.
4 Experiments
4.1 Experimental Setup
@@
Training. We use LSDIR [14] as the training data with the $512\times 512$ patch size. To generate paired HQ-LQ training data, we follow the degradation pipeline from Real-ESRGAN [25]. We use AdamW optimizer [19] with a learning rate $5\times 10^{-5}$ and set LoRA rank to 4 for both the student model and LoRA model. We employ the SD 2.1-base as the pre-trained model and fine-tune it for 2k iterations using 8 NVIDIA A40 GPUs with a batch size of 24. For the choice of hyperparameters, we set $\lambda=0.5$ and $\gamma=0$ in Eq. (5). We adopt the degradation-aware prompt extraction (DAPE) module [33] to extract text prompts.
@@
Test Dataset. We evaluate our method in both synthetic and real-world dataset. For the synthetic dataset, we randomly crop 3K patches with a resolution of 512 × 512 from the DIV2K [1] validation set and synthesize LQ data using the same pipeline as that in training. For real-world data, we employ RealSR [2], DrealSR [30], and RealLR200 [33]. We center-crop RealSR [2] and DrealSR [30] datasets with size $128\times 128$ for LQ images and $512\times 512$ for HQ image. For RealLR200 [33] dataset, since the corresponding HQ images are unavailable, we perform only a $128\times 128$ center-crop on the LQ images.
@@
Evaluation Metrics. We utilize several reference and non-reference metrics to evaluate the performance of various methods on the test data. For the reference measures, we employ PSNR, SSIM [29], and LPIPS [41] to measure image fidelity. For the non-reference measures, we employ CLIPIQA [23], MUSIQ ([13], MANIQA [36], TOPIQ [4], and QALIGN [31] to measure image quality.
@@
Compared Methods. We compare our method with several multi-step diffusion-based methods StableSR [24], DiffBIR [18], SeeSR [33], and one-step methods SinSR [26], OSEDiff [32], S3Diff [39], AdcSR [3], TSDSR [8], and PisaSR [21]. All comparative results are obtained using publicly released code and model weights for testing.
4.2 Comparisons with State-of-the-art Methods
@@
Quantitative Comparisons. We set up the timestep condition $t_{s}=500$ in our method, and show the quantitative comparisons on the four synthetic and real-world datasets in Table 1. We have the following observations: (1) TADSR achieves the highest no-reference scores across four datasets, except for the MUSIQ on DIV2K-Val. This demonstrates that TADSR can more effectively leverage the generative priors from SD to produce more realistic results. Notably, TADSR is the only one-step method that consistently outperforms multi-step methods on all no-reference metrics, achieving both efficiency and perceptual quality. (2) TADSR maintains PSNR values comparable to other SD-based one-step methods, indicating a good balance between fidelity and realism. (3) TADSR shows clear improvements over other SD-based one-step methods on CLIPIQA and TOPIQ, highlighting its superior semantic awareness and generative capability. (4) Under the same number of parameters as OSEDiff, TADSR significantly outperforms OSEDiff on no-reference metrics while maintaining comparable reference metrics, further demonstrating TADSR’s effective utilization of generative priors and the resulting performance improvements.
@@
Qualitative Comparisons. Figure 5 shows the visual comparisons between our method and the other state-of-the-art Real-ISR methods. As shown in the first row, TADSR generates significantly more natural and sharper textures from heavily degraded LQ images, especially in facial regions such as the teeth, eyes, and eyebrows, demonstrating its strong semantic generation capability. In the second row, the digits and letters produced by TADSR appear much clearer, showcasing its superior degradation removal ability while preserving fidelity. In the third row, TADSR yields more natural results around the eagle’s eyes and beak. In the fourth row, only TADSR accurately restored natural-looking facial features such as the nose, mouth, and chin. Other methods generally suffered from degradation, resulting in some distortion, and failed to reconstruct a plausible chin structure. Overall, thanks to the ability to distill generative priors from SD more effectively in TAVSD loss, TADSR can produce natural and realistic results in a single diffusion step. Compared to other methods, it achieves strong perceptual quality while maintaining high efficiency.
4.3 Ablation Study
Figure 6: (a) Quantitative metrics of our method under different timestep $t_{s}$, evaluated on the DrealSR dataset. (b) Comparison of our method under different timesteps $t_{s}$, PisaSR under different semantic guidance weights $\lambda_{sem}$, and other one-step diffusion-based Real-ISR methods, evaluated on the DrealSR dataset.
@@ 4.3 Ablation Study
Impact of Different Timestep Condition. As shown in Figure 6(a), we analyze the impact of timestep $t_{s}$ in our method on both reference and no-reference metrics. As $t_{s}$ increases, PSNR exhibits a decreasing trend while QALIGN shows an upward trend, indicating a trade-off where fidelity is sacrificed to enhance realism. This trade-off between fidelity and realism aligns with the function of $t_{s}$, as a larger $t_{s}$ means that TAVSD provides stronger generative guidance, while a smaller $t_{s}$ provides more fidelity-preserving guidance. Similar visual results can be observed in supplementary materials. Furthermore, we compare the results of our method under different $t_{s}$, PisaSR under different $\lambda_{sem}$ settings, and other one-step Real-ISR methods, as shown in Figure 6(b). It can be observed that our method consistently lies in the top-right corner across different $t_{s}$. When $t_{s}$ equals 200, our method achieves 26.61dB PSNR, which is more than 1dB higher than SinSR, and QALIGN is significantly higher than SinSR. In contrast, although PisaSR can also achieve a PSNR of 29.60dB by tuning the $\lambda_{pix}=1.0$ and $\lambda_{sem}=0.6$, its QALIGN is only 2.91, which is similar to SinSR. This indicates that our method achieves a substantial improvement in fidelity with only a minimal compromise in realism.
@@
Figure 7: Vision Comparisons of the ablation study on TAE and TAVSD. Baseline use the original VAE encoder and VSD loss.
@@
Table 2: Quantitative Comparison of ablation study on TAVSD and TAE. Baseline uses the original VAE encoder and VSD loss.
@@
Impact of TAVSD and TAE. To validate the effectiveness of TAVSD and TAE, we conducted ablation studies by removing them. We employ the original VAE encoder in SD and the VSD loss as our baseline and conduct ablation studies by separately removing TAE and TAVSD. We use PSNR to evaluate fidelity and CLIPIQA, MUSIQ, and TOPIQ to assess realism. As shown in Table 2, we have the following three key observations: (1) Although the baseline also adopts randomly sampled timesteps during training, after removing TAE, both reference and no-reference metrics decline, demonstrating that timestep-adaptive latent distribution plays a crucial role in effectively utilizing the generative priors in SD. (2) When TAVSD is ablated, all metrics similarly decrease, indicating that more consistent guidance from the teacher model better activates generative priors across different timesteps. (3) Baseline shows significant degradation in PSNR and moderate decline in others, proving that both TAE and TAVSD improve fidelity and realism. Additionally, Figure 7 presents a visual comparison of our ablation studies, showing that both the absence of TAE/TAVSD leads to unrealistic parrot reconstructions, while the baseline even produces visible artifacts. In contrast, our method produces realistic and natural results by fully exploiting the generative priors in SD.
5 Conclusion